In one form of plethysmograph (a device for measuring volume), a rubber capillary tube with an inside diameter of 1.39 mm is filled with mercury at 20
Answer #1

(a) R= L/A

therefor R=9.4 *10^-7 * 1/ r^2

R= 9.4 *10^-7 / 1.51 * 10^-6

therefor R =0.622 ohm

(b) V = Ai/Li = Af/Lf and Af = Ai(Li/Lf).

Af =1.508 * 10^-6

therefore Fractional change in resistanve = new resistance- initial resistance

new resistance=0.6239 ohm

Fractional change= 0.6239 ohm -0.622 ohm

=0.0019 ohm